# The Design Space of Lockfiles Across Package Managers

### v2

@@ 7 Conclusion
We have conducted the first in-depth study of lockfiles across 7 package managers. We analyzed their documentation and implementation, and collected quantitative data from open-source GitHub projects and qualitative data from developer interviews. We systematically documented how different package managers follow different approaches to generate and enforce lockfiles, and these decisions ultimately result in varied levels of developer satisfaction. We further identified the reasons why developers find lockfiles useful: to achieve deterministic builds, to perform integrity checks, to simplify debugging. We also unveiled the unique challenges developers face when using lockfiles: readability issues, indirect dependency locking limitations, and delayed dependency updates. Based on these observations, we provided 2 recommendations for package manager users and 3 recommendations for package manager developers.
Today, only the Go ecosystem systematically generates and adds lockfiles in version control system. The Go lockfile is particularly readable, concise and clearly separates the list of dependencies from their checksums, making them easier to understand for developers. On the other hand, lockfiles are almost absent from the Java ecosystem. Out of the 7 package managers we study, only Gradle does not generate a lockfile by default. Meanwhile, Maven, the other major package manager for Java does not have a lockfile at all. We recommend the Maven community to add this feature and learn from the best practices to design an informative and usable lockfile.
@@ Acknowledgments
We thank all the interview participants for dedicating their time for this work. We acknowledge the valuable insights provided by Ruy Adorno, a software engineer at vlt. We also thank Ian Arawjo for feedback on early versions of this paper. This work is supported by the CHAINS project funded by Swedish Foundation for Strategic Research (SSF), by the Wallenberg Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation, and by IVADO and the Canada First Research Excellence Fund.

### v3

@@ 8 Conclusion
We have conducted the first in-depth study of lockfiles across seven package managers. We analyzed their documentation and implementation, and collected quantitative data from public GitHub projects and qualitative data from developer interviews. We systematically documented how different package managers follow different approaches to generate and enforce lockfiles, and these decisions ultimately result in varied levels of developer satisfaction. We further identified the reasons why developers find lockfiles useful: to achieve deterministic builds, to perform integrity checks, to simplify debugging. We also unveiled the unique challenges developers face when using lockfiles: readability issues, indirect dependency locking limitations, and delayed dependency updates. Based on these observations, we provided two recommendations for package manager users and three recommendations for package manager developers.
Today, only the Go ecosystem systematically generates and adds lockfiles in version control system. The Go lockfile is particularly readable, concise and clearly separates the list of dependencies from their checksums, making them easier to understand for developers. On the other hand, lockfiles are almost absent from the Java ecosystem. Out of the seven package managers we study, only Gradle does not generate a lockfile by default. Meanwhile, Maven, the other major package manager for Java does not have a lockfile at all. We recommend the Maven community to add this feature and learn from the best practices to design an informative and usable lockfile.
@@ Acknowledgments
We thank all the interview participants for dedicating their time for this work. We acknowledge the valuable insights provided by Ruy Adorno, a software engineer at vlt. We also thank Ian Arawjo for feedback on the interview protocol and early versions of this paper.
Declarations
Funding
This work is supported by the CHAINS project funded by Swedish Foundation for Strategic Research (SSF), by the Wallenberg Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation, and by IVADO and the Canada First Research Excellence Fund.
Data Availability Statement
All materials related to the qualitative analysis that cannot be traced back to the interview participants, including interview invitation emails, interview protocol, the codebook, as well as all data from the quantitative analysis, are available at https://github.com/chains-project/lockfiles-comprehensive-study.
Ethical Considerations
This paper includes minimal-risk data collected from developers during interviews. Participation in the study was entirely voluntary. All data collected from participants were anonymized and were handled in accordance with the Privacy Act. Prior to the study, we obtained approval from our organization’s Research Ethics Board (REB).